Sunspot Beverages, Ltd., of Fiji uses the weighted-average method in its process costing system. It makes blended tropical fruit drinks in two stages. Fruit juices are extracted from fresh fruits and then blended in the Blending Department. The blended juices are then bottled and packed for shipping in the Bottling Department. The following information pertains to the operations of the Blending Department for June.

Percent Completed

Units

Materials

Conversion

Work in process, beginning

20,000

100%

75%

Started into production

180,000

Completed and transferred out

160,000

Work in process, ending

40,000

100%

25%

Materials

Conversion

Work in process, beginning

$25,200

$24,800

Cost added during June

$334,800

$238,700

1. Calculate the Blending Department's equivalent units of production for materials and conversion in June.

1.

Equivalent units of production:

Materials

Conversion

Transferred to next department........................

Equivalent units in ending work in process inventory:

Materials: 40,000 units x 100% complete.......

Conversion: 40,000 units x 25% complete......

Equivalent units of production...........................

2. Calculate the Blending Department's cost per equivalent unit for materials and conversion in June.

2.

Cost per equivalent unit

Materials

Conversion

Cost of beginning work in process.................

Cost added during the period.........................

Total cost (a).................................................

Equivalent units of production (b)..................

Cost per equivalent unit (a) ÷ (b)..................

3. Calculate the Blending Department's cost of ending work in process inventory for materials, conversion, and in total for June. Also calculate the Blending Department's cost of units transferred out to the Bottling Department for materials, conversion, and in total for June.

Assigning costs to units:

Materials

Conversion

Total

Ending work in process inventory:

Equivalent units....................

Cost per equivalent unit........

Cost of ending work in process inventory.............................

Units completed and transferred out:

Units transferred to the next department.........................

Cost per equivalent unit........

Cost of units completed and transferred out....................

4. Prepare a cost reconciliation report for the Blending Department for June.

Cost reconciliation:

Costs to be accounted for:

Cost of beginning work in process inventory

Costs added to production during the period

Total cost to be accounted for........................

Costs accounted for as follows:

Cost of ending work in process inventory.......

Cost of units completed and transferred out..

Total cost accounted for.................................
Answer #1

B C D Materials 160,000 Conversion 160,000 1 1) Equivalent units of production: 2 3 Transferred to next department 4 Equivale28 4) Cost reconciliation: 29 30 Costs to be accounted for: 31 Cost of beginning work in process inventory 32 Costs added to

B C D Materials =160000*100% Conversion =160000*100% 1 1) Equivalent units of production: 2 3 Transferred to next department27 =25200+24800 =334800+238700 28 4) Cost reconciliation: 29 30 Costs to be accounted for: 31 Cost of beginning work in proce